# Towards Secure Virtual Elections: Multiparty Computation of Order Based Voting Rules

### v4

@@ Abstract
Electronic voting systems have significant advantages in comparison with physical voting systems. One of the main challenges in e-voting systems is to secure the voting process: namely, to certify that the computed results are consistent with the cast ballots and that the voters’ privacy is preserved. We propose herein a secure voting protocol for elections that are governed by order-based voting rules. Our protocol offers perfect ballot secrecy in the sense that it issues only the required output while no other information on the cast ballots is revealed. Such perfect secrecy, achieved by employing secure multiparty computation tools, may increase the voters’ confidence and, consequently, encourage them to vote according to their true preferences. Evaluation of the protocol’s computational costs establishes that it is lightweight and can be readily implemented in real-life electronic elections. Keywords: Secure voting, Perfect ballot secrecy, Multiparty computation, Computational social choice, Virtual elections
@@ 1 Introduction
Electronic voting has significant advantages in comparison with the more commonplace physical voting: it is faster, it reduces costs, it is more sustainable, and in addition, it may increase voters’ participation. One of the main challenges in holding e-voting is to secure the voting process: namely, to make sure that it is secure against attempted manipulations so that the computed results are consistent with the cast ballots and that the privacy of the voters is preserved.
The usual meaning of voter privacy is that the voters remain anonymous. Namely, the linkage between ballots and the voters that cast them remains hidden, thus preserving anonymity. However, the final election results and the full tally (i.e., all ballots) are revealed. While such information exposure may be deemed benign or even desired, in some cases, it may be problematic. For example, in small-scale elections (say, elections to university senates), the full tally may put the voters’ privacy at risk; say, if Alice promised Bob to vote for him, but at the end of the day, no one had voted for him, Alice’s failure to keep her promise to Bob would be exposed. Such potential privacy breaches may be an obstacle for some voters to vote truthfully. In some cases, exposing the full tally might severely damage the image of some candidates who may, consequently, refrain from submitting their candidacy again in future elections. Such problems can be averted if the voting system is tally-hiding (Szepieniec and Preneel, 2015); namely, if it reveals only the desired final results and nothing else. The final results could be just the identity of the winner (say, when selecting an editor-in-chief or a prime minister), $K>1$ winners (e.g., when needing to select $K$ new board members), $K>1$ winners with their ranking (say, when needing to award first, second and third prizes), or a score for each candidate (as is the case in elections for parliament, where the candidates are parties and the score for each party is the number of seats in the parliament).
@@
In this paper, we design a tally-hiding voting system that offers perfect ballot secrecy, or full privacy (Chaum, 1988); i.e., given any coalition of voters, the protocol does not reveal any information on the ballots beyond what can be inferred from the published results. Such full privacy may reduce the possibility of voters’ coercion and also increase the voters’ confidence that their vote remains secret. Hence, full privacy may encourage voters to vote according to their true preferences.
@@
There are various families of voting rules that can be used in elections. For example, in score-based voting rules, a score for each candidate is computed subject to the specifications of the underlying voting rule, and then the winning candidate is the one whose aggregated score is the highest. In this study we focus on order-based (a.k.a pairwise-comparison) voting rules, where the relative order of the candidates is considered by the underlying voting rule (Brandt and Sandholm, 2005).
@@
In order to provide privacy for the voters, it is necessary to protect their private data (i.e. their ballots) from the tallier while still allowing the tallier to perform the needed computations on the ballots in order to output the required election results. The cryptographic tool that is commonly used towards that end is homomorphic encryption, namely, a form of encryption that preserves algebraic structure and thus enables the performance of meaningful computations on the ciphertexts. However, homomorphic encryption has a significant computational overhead. The main idea that underlies our suggested system is to use distributed tallying. Namely, our system involves $D>1$ independent talliers to whom the voters send information relating to their private ballots. With such distributed tallying, it is possible to replace the costly cryptographic protection shield of homomorphic encryption with the much lighter-weight cryptographic shield of secret sharing. The talliers then engage in specially designed protocols of multiparty computation that allow them to validate that each cast ballot is a legitimate ballot (in the sense that it complies with the specifications of the underlying voting rule) and then to compute from the cast ballots the required election results, while still remaining totally oblivious to the content of those ballots.
A high-level description of our system is presented in Figure 1. The first phase is the voting phase: the voters send to each of the talliers messages (shares) relating to their secret ballots. In the second phase, the talliers communicate amongst themselves in order to validate each of the incoming ballots (which remain secret to them) and then, at the end of the day, perform the tallying over all legal ballots according to the underlying rule. Finally, the talliers broadcast the results back to the voters.
@@
Our contributions. We consider elections that are governed by an order-based voting rule. We focus on two such rules – Copeland (Copeland, 1951), and Maximin (Simpson, 1969) (a.k.a Kramer-Simpson). We devise a fully private protocol for computing the election results. The election output is a ranking of the candidates from which the winning candidate(s) can be determined. Our protocol is lightweight and can be readily implemented in virtual elections. In Appendices G and H we describe two other order-based rules, Kemeny-Young (Kemeny, 1959; Young, 1995) and Modal Ranking (Caragiannis et al., 2014), and describe the extension of our methods for those rules as well.
The paper is structured as follows. In Section 2 we explain the cryptographic foundations on which our protocol is based. In Section 3 we present our secure voting protocol for Copeland and Maximin rules and then, in Section 4, we evaluate its computational costs. A survey of related work is given in Section 5, and we conclude in Section 6.
2 Cryptographic preliminaries
Our protocol relies heavily on cryptographic machinery. This section provides the required cryptographic background and an explanation of how the presented techniques are implemented in our protocol. In Section 2.1 we provide a brief introduction to secret sharing. In Section 2.2 we describe the cryptographic principles of our protocol. Then, in the following sections (Sections 2.3-2.6) we provide the details of specific computations that we use in our protocol.
We note that while Sections 2.1 and 2.2 are necessary for understanding our protocol that is presented in Section 3, they are also sufficient, in the sense that readers who are less interested in the cryptographic details may skip Sections 2.3-2.6 and still be able to fully understand the secure voting protocol.
@@ 2.1 Secret sharing
Secret sharing schemes (Shamir, 1979) enable distributing a secret $s$ among a set of parties, $\mbox{\bf T}=\{T_{1},\ldots,T_{D}\}$. Each party, $T_{d}$, $d\in[D]:=\{1,\ldots,D\}$, is given a random value $s_{d}$, called a share, that relates to the secret $s$. Those shares satisfy the following conditions: (a) $s$ can be reconstructed only by combining the shares given to specific subsets of parties, which are called authorized subsets; (b) shares held by any other subset of parties reveal zero information on $s$.
The notion of secret sharing was introduced independently, by Shamir (1979) and Blakley (1979), for the case of threshold secret sharing. In threshold secret sharing there is some threshold $D^{\prime}\leq D$ and then the authorized subsets are those of size at least $D^{\prime}$. Such secret sharing schemes are called $D^{\prime}$-out-of-$D$.
@@
$\bullet$ ${\sf Reconstruct}_{D^{\prime}}(s_{1},\ldots,s_{D})$. The procedure is given any selection of $D^{\prime}$ shares out of $\{s_{1},\ldots,s_{D}\}$; it then interpolates a polynomial $g(\cdot)$ of degree at most $D^{\prime}-1$ using the given points and outputs $s=g(0)$. Clearly, any selection of $D^{\prime}$ shares out of the $D$ shares will yield the same polynomial $g(\cdot)$ that was used to generate the shares, as $D^{\prime}$ point values determine a unique polynomial of degree at most $D^{\prime}-1$. Hence, any selection of $D^{\prime}$ shares will issue the secret $s$. On the other hand, any selection of $D^{\prime}-1$ shares reveals nothing about the secret.
2.2 The cryptographic principles of our protocol
Our protocol involves two sets of parties: voters, $\mbox{\bf V}=\{V_{1},\ldots,V_{N}\}$, and talliers, $\{T_{1},\ldots,T_{D}\}$. The talliers are assumed to be semi-honest, i.e., they follow the prescribed protocol, but try to extract from their view in the protocol information on the private inputs of the voters. We also assume them to have an honest majority, in the sense that if some of them are corrupted by a malicious adversary, their number is smaller than
$D^{\prime}:=\lfloor(D+1)/2\rfloor\,.$ (1)
The voters, on the other hand, may be malicious. Hence, our protocol includes mechanisms to detect potential deviations from the prescribed protocol in order to ensure that the computed election results correctly identify the winners as dictated by all legal ballot and the underlying rule. During the entire process, the sensitive information that should be kept secret are the private ballots as well as all intermediate computed values; nothing beyond the identity of the winner(s) should be leaked.
@@ 2.2 The cryptographic principles of our protocol
In the protocol, each voter creates shares of his [^1] For the sake of simplicity, we keep referring to parties by the pronoun “he”. In our context, those parties may be voters, who are humans of any gender, or talliers, that are typically (genderless) servers. private ballot and distributes them to the $D$ talliers. As those ballots are matrices (see the definitions in Section 3), the secret sharing is carried out for each matrix entry independently.
@@
We use Shamir’s $D^{\prime}$-out-of-$D$ secret sharing, where hereinafter $D^{\prime}$ is as defined in Eq. (1). With that selection of the threshold $D^{\prime}$, at least half of the talliers would need to collude in order to recover the secret ballots. Under our assumption that the set of talliers has an honest majority, our protocol provides information-theoretic security. Higher values of $D$ will imply greater security against coalitions of corrupted talliers, but they will also imply higher costs.
@@
The first thing that the talliers need to do when receiving from a voter his ballot matrix shares is to verify that those shares correspond to a legal ballot (see Section 3.2 where we explain what constitutes a legal ballot). Then, at the end of the election period, the talliers need to compute the identity of the winning candidates, as dictated by the rule. Those tasks would be easy if the talliers could use their shares in order to recover the ballots. However, they must not do so, in order to protect the voters’ privacy. Instead, they must perform those computations on the distributed shares, without revealing the shares to each other. As we shall see later on, those computations boil down to the four specific tasks that we proceed to describe.
@@
Let $x_{1},\ldots,x_{L}$ be $L$ secrets in the underlying field ${\bf Z}_{p}$ ($L$ is any integer) and assume that the talliers hold $D^{\prime}$-out-of-$D$ shares in each one of those secrets. Then the talliers need to perform the following computational tasks without recovering the secrets $x_{1},\ldots,x_{L}$ or learn anything about them:
@@
    If $y=f(x_{1},\ldots,x_{L})$, where $f$ is some public arithmetic function, compute $D^{\prime}$-out-of-$D$ shares in $y$.
@@
    If $y=1_{x_{1}<x_{2}}$ is the bit that equals 1 if $x_{1}<x_{2}$ and 0 otherwise [^2] Hereinafter, for any predicate $\Pi$, we let $1_{\Pi}$ denote the binary variable that equals 1 iff (if and only if) the predicate $\Pi$ holds., compute $D^{\prime}$-out-of-$D$ shares in $y$.
@@
    If $p>2N$ ($N$ is the number of voters) and $x_{1}\in[-N,N]$, compute $D^{\prime}$-out-of-$D$ shares in the bit $1_{x_{1}>0}$.
@@
    If $p>2N$ and $x_{1}\in[-N,N]$, compute $D^{\prime}$-out-of-$D$ shares in the bit $1_{x_{1}=0}$.
In the next subsections, we describe sub-protocols of secure multiparty computation (MPC) (Yao, 1982) for performing those computations.
2.3 Evaluating polynomials
Let $f(x_{1},\ldots,x_{L})$ be a polynomial in $L$ variables. Assume that the talliers have $D^{\prime}$-out-of-$D$ shares in each of the $L$ inputs and they wish to compute $D^{\prime}$-out-of-$D$ shares in the output $y=f(x_{1},\ldots,x_{L})$ To do so, it is necessary to solve the following problems: given $D^{\prime}$-out-of-$D$ shares in two secret values $u,v\in{\bf Z}_{p}$, compute $D^{\prime}$-out-of-$D$ shares in $au+bv$, where $a,b$ are public field elements, and in $u\cdot v$, without reconstructing any of those values ($u$, $v$, $au+bv$ and $u\cdot v$).
Let $u_{d}$ and $v_{d}$ denote the shares held by $T_{d}$, $d\in[D]$, in the two input values $u$ and $v$, respectively. The linearity of secret sharing implies that $au_{d}+bv_{d}$, $d\in[D]$, are proper $D^{\prime}$-out-of-$D$ shares in $au+bv$. Hence, linear combinations may be resolved without any interaction amongst the talliers.
The procedure for resolving multiplication is trickier and requires the talliers to interact. In our protocol, we use the construction proposed by Damgård and Nielsen (2007), enhanced by a work by Chida et al. (2018) that demonstrates some performance optimizations. The details of this computation are not essential for understanding our secure voting protocol, but for the sake of completeness we describe the computation in Appendix A.
The complexity of securely evaluating arithmetic functions is measured by the number of multiplications (as those are the operations that require significantly more computational and communication costs), and the degree of the polynomial (as it determines the number of rounds of communication).
2.4 Secure Comparison
Assume that $T_{1},\ldots,T_{D}$ hold $D^{\prime}$-out-of-$D$ shares in two integers $a$ and $b$, where both $a$ and $b$ are smaller than $p$, which is the size of the underlying field ${\bf Z}_{p}$. They wish to compute $D^{\prime}$-out-of-$D$ secret shares in the bit $1_{a<b}$ without learning any other information on $a$ and $b$. A protocol that does that is called secure comparison.
Nishide and Ohta (2007) presented a method for secure comparison that is based on the following simple observation. Let us denote the bits $1_{a<\frac{p}{2}}$, $1_{b<\frac{p}{2}}$, $1_{[(a-b)\mod p]<\frac{p}{2}}$, and $1_{a<b}$ by $w$, $x$, $y$, $z$, respectively. Then
$z$ $=w(1-x)+(1-w)(1-x)(1-y)+wx(1-y)$ (2)
$=1-x-y+xy+w(x+y-2xy)\,.$
Hence, we reduced the problem of comparing two secret shared values, $a$ and $b$, to computing three other comparison bits — $w,x,y$, and then evaluating an arithmetic function of them, Eq. (2). What makes this alternative expression efficiently computable is the fact that in the three comparison bits, $w=1_{a<\frac{p}{2}}$, $x=1_{b<\frac{p}{2}}$, and $y=1_{[(a-b)\mod p]<\frac{p}{2}}$, the right-hand side is $\frac{p}{2}$, as we proceed to explain.
Lemma 1.
Given a finite field ${\bf Z}_{p}$ and a field element $q\in{\bf Z}_{p}$, then $q<\frac{p}{2}$ iff the least significant bit (LSB) of ($2q$ mod $p$) is zero.
(The proofs of all lemmas and theorems are given in Appendices BF.)
In view of Lemma 1, the talliers may compute $D^{\prime}$-out-of-$D$ shares in $w=1_{a<\frac{p}{2}}$ (and similarly for $x=1_{b<\frac{p}{2}}$, and $y=1_{[(a-b)\mod p]<\frac{p}{2}}$) as follows: each $T_{d}$ will translate the share he holds in $a$ to a share in $2a$ by multiplying the share by 2; then, all talliers will use their shares in $2a$ in order to compute shares in the LSB of $2a$ Nishide and Ohta (2007).
We conclude this section by commenting on the complexity of the above described secure comparison protocol. Computing shares in the LSB of a shared value requires $13$ rounds of communication and $193\ell+$ 1 multiplications, where hereinafter $\ell=\log_{2}(p)$. Since we have to compute three such bits (i.e., $w$, $x$, and $y$) then we can compute shares of those three bits in 13 rounds and a total of $3279\ell+$ 3 multiplications. Finally, we should evaluate the expression in Eq. (2), which entails two additional rounds and two additional multiplications. Hence, the total complexity is $15$ rounds and $5279\ell+$ 5 multiplications.
2.5 Secure testing of positivity
Let $x$ be an integer in the range $[-N,N]$. Assume that $T_{1},\ldots,T_{D}$ hold $D^{\prime}$-out-of-$D$ shares in $x$, where the underlying field is ${\bf Z}_{p}$, and $p>2N$. Our goal here is to design an MPC protocol that will issue to $T_{1},\ldots,T_{D}$ $D^{\prime}$-out-of-$D$ shares in the bit $1_{x>0}$, without learning any further information on $x$.
Lemma 2.
Under the above assumptions, $x>0$ iff the LSB of $(-2x\mod p)$ is 1.
Lemma 2 implies that testing the positivity of a secret requires a single LSB computation. Hence, in view of the discussion in Section 2.4, the computational cost of that task is $13$ rounds of communication and $193\ell+$ 1 multiplications.
2.6 Secure testing of equality to zero
Let $x$ be an integer in the range $[-N,N]$. Assume that $T_{1},\ldots,T_{D}$ hold $D^{\prime}$-out-of-$D$ shares in $x$, where the underlying field is ${\bf Z}_{p}$, and $p>2N$. Our goal here is to design an MPC protocol that will issue to $T_{1},\ldots,T_{D}$ $D^{\prime}$-out-of-$D$ shares in the bit $1_{x=0}$, without learning any further information on $x$. It is possible to solve that problem by implementing the MPC positivity testing that we described above in Section 2.5, once for $x$ and once for $-x$. Clearly, $x=0$ iff both of those tests fail. However, we can solve that problem in a much more efficient manner, as we proceed to describe.
Fermat’s little theorem states that if $x\in{\bf Z}_{p}\setminus\{0\}$ then $x^{p-1}=1$ mod $p$. Hence, $1_{x\neq 0}=\left(x^{p-1}\mod p\right)$. Therefore, shares in the bit $1_{x\neq 0}$ can be obtained by simply computing $x^{p-1}$ mod $p$. The latter computation can be carried out by the square-and-multiply algorithm with up to $2\ell$ consecutive multiplications, where, as before, $\ell=\log p$. Finally, as $1_{x=0}=1-1_{x\neq 0}$, then shares in $1_{x\neq 0}$ can be readily translated into shares in $1_{x=0}$. The cost of the above described computation is significantly smaller than the cost of the alternative approach that performs positivity testing of both $x$ and $-x$.
3 A secure order based voting protocol

### v7

# Secure Order Based Voting Using Distributed Tallying

@@ Abstract
Electronic voting systems have significant advantages in comparison with physical voting systems. One of the main challenges in e-voting systems is to secure the voting process: namely, to certify that the computed results are consistent with the cast ballots and that the voters’ privacy is preserved. We propose herein a secure voting protocol for elections that are governed by order-based voting rules. Our protocol, in which the tallying task is distributed among several independent talliers, offers perfect ballot secrecy in the sense that it issues only the required output while no other information on the cast ballots is revealed. Such perfect secrecy, achieved by employing secure multiparty computation tools, may increase the voters’ confidence and, consequently, encourage them to vote according to their true preferences. We implemented a demo of a voting system that is based on our protocol and we describe herein the system’s components and its operation. Our implementation demonstrates that our secure order-based voting protocol can be readily implemented in real-life large-scale electronic elections.
@@ 1 Introduction
Electronic voting has significant advantages in comparison with the more commonplace physical voting: it is faster, it reduces costs, it is more sustainable, and in addition, it may increase voters’ participation. In recent years, due to the COVID-19 pandemic and the need for social distancing, e-voting platforms have become even more essential. One of the main challenges in holding e-voting is to secure the voting process: namely, to make sure that it is secure against attempted manipulations so that the computed results are consistent with the cast ballots and that the privacy of the voters is preserved.
The usual meaning of voter privacy is that the voters remain anonymous. Namely, the linkage between ballots and the voters that cast them remains hidden, thus preserving anonymity. However, the final election results and the full tally (i.e., all ballots) are revealed. While such information exposure may sometimes be considered benign or even desired, in some cases it may be problematic. For example, in small-scale elections—such as those for a university senate—publishing the full tally may compromise ballot secrecy. Suppose Bob is a candidate and only a small number of voters participate. If the final tally reveals that Bob received zero votes, and it is known or assumed that Alice supported him, observers (including Bob himself) may infer that Alice did not follow through. Such inference risks may deter voters from expressing their true preferences, either out of fear of social consequences or due to pressure from candidates or peers, particularly in close-knit communities.
In contrast to weaker privacy notions, such as anonymity or partial tally hiding, perfect ballot secrecy ensures that nothing is disclosed except for the final results. The final results could be just the identity of the winner (say, when selecting an editor-in-chief or a prime minister), $K>1$ winners (e.g., when needing to select $K$ new board members), $K>1$ winners with their ranking (say, when needing to award first, second and third prizes), or a score for each candidate (as is the case in elections for parliament, where the candidates are parties and the score for each party is the number of seats in the parliament).
@@
In this paper, we present the design of a tally-hiding voting system that ensures perfect ballot secrecy—a concept sometimes referred to as full privacy (see, e.g., [13]). Such secrecy ensures that given any coalition of voters, the protocol does not reveal any information on the ballots beyond what can be inferred from the published results. Such perfect ballot secrecy may reduce the possibility of voters’ coercion and increase their confidence that their vote remains secret. Hence, it may encourage voters to vote according to their true preferences.
@@
There are various families of voting rules that can be used in elections. For example, in score-based voting rules, a score for each candidate is computed subject to the specifications of the underlying voting rule, and then the winning candidate is the one whose aggregated score is the highest. In this study, we focus on order-based (a.k.a pairwise-comparison) voting rules, where the relative order of the candidates is considered by the underlying voting rule [9].
Our contributions. We consider a scenario in which there is a set of voters that hold an election over a given set of candidates, where the election is governed by an order-based voting rule. We devise a fully private protocol for computing the election results. The election output is a ranking of the candidates from which the winning candidate(s) can be determined. Our protocol is lightweight and can be readily implemented in virtual elections.
@@
In order to provide privacy for the voters, it is necessary to protect their private data (i.e. their ballots) from the tallier while still allowing the tallier to perform the needed computations on the ballots in order to output the required election results. The cryptographic tool that is commonly used towards this end is homomorphic encryption, namely, a form of encryption that preserves the algebraic structure and thus enables the performance of meaningful computations on the ciphertexts. However, homomorphic encryption has a significant computational overhead. The main idea that underlies our suggested system is to use distributed tallying. Namely, our system involves $D>1$ independent talliers to whom the voters send information relating to their private ballots. With such distributed tallying, it is possible to replace the costly cryptographic protection shield of homomorphic encryption with the much lighter-weight cryptographic shield of secret sharing (all relevant cryptographic background is provided in Section 2). The talliers then engage in specially designed protocols of multiparty computation that allow them to validate that each cast ballot is a legitimate ballot (in the sense that it complies with the specifications of the underlying voting rule) and then to compute from the cast ballots the required election results, while still remaining totally oblivious to the content of those ballots. A high-level description of our system is presented in Figure 1. The first phase is the voting phase: the voters send to each of the talliers messages (shares) relating to their secret ballots. In the second phase, the talliers communicate among themselves in order to validate each of the incoming ballots (which remain secret to them) and then, at the end of the day, perform the tallying over all legal ballots according to the underlying rule. Finally, the talliers broadcast the results back to the voters.
@@
The bulk of this study is devoted to two order-based voting rules – Copeland [16], and Maximin [48] (a.k.a Kramer-Simpson), which are formally defined in Section 3.1. In Appendix B we describe two other rules in this family, Kemeny-Young [32, 52] and Modal Ranking [11], and describe the extension of our methods for those rules as well.
The paper is structured as follows. In Section 2 we explain the cryptographic principles of distributed tallying. In Section 3 we present our secure voting protocols for Copeland and Maximin rules. We focus there on the case where voters must determine a strict ranking of the candidates. Then, in Section 4, we extend our discussion to the more involved case in which the voters’ rankings may include ties between some of the candidates. In Section 5 we describe an implementation of a secure voting system that is based on our protocols; our implementation is in Python and is open source. In Section 6 we survey related work. Lastly, Section 7 provides concluding comments, as well as a description of an implementation of our presented protocol and system in the Gentoo [^1] https://www.gentoo.org/ council elections.
The Appendix includes additional discussions. In A we describe cryptographic sub-protocols that we use in our protocol, while in B we discuss secure protocols for two other order-based voting rules -- Kemeny-Young and Modal Ranking.
2 Distributed tallying
A key principle of our protocol is distributing the tallying task among multiple independent talliers. This distribution is supported by secret sharing [45] and secure multiparty computation (MPC) [51]. In Section 2.1 we give a brief recap of secret sharing. Then, in Section 2.2 we describe the distributed tallying on which our protocol is based.
Our protocol utilizes MPC sub-protocols that perform secure computations over secret shares. We treat those sub-protocols as black boxes. We describe their inputs and outputs in Section 2.2. Readers seeking a deeper understanding of these MPC components may refer to A.
@@ 2.1 Secret sharing
Secret sharing schemes [45] enable distributing a secret $s$ among a set of parties, $\mbox{\bf T}=\{T_{1},\ldots,T_{D}\}$. Each party, $T_{d}$, $d\in[D]:=\{1,\ldots,D\}$, is given a random value $s_{d}$, called a share, that relates to the secret $s$. Those shares satisfy the following two conditions: (a) $s$ can be reconstructed only by combining the shares given to specific authorized subsets of parties, and (b) shares belonging to unauthorized subsets of parties reveal zero information on $s$. In threshold secret sharing there is some threshold $D^{\prime}\leq D$ and then the authorized subsets are those of size at least $D^{\prime}$. Such secret sharing schemes are called $D^{\prime}$-out-of-$D$.
@@
$\bullet$ ${\sf Reconstruct}_{D^{\prime}}(s_{1},\ldots,s_{D})$. The procedure is given any selection of $D^{\prime}$ shares out of $\{s_{1},\ldots,s_{D}\}$; it then interpolates a polynomial $g(\cdot)$ of degree at most $D^{\prime}-1$ using the given point values, and it outputs $s=g(0)$. Clearly, any selection of $D^{\prime}$ shares out of the $D$ shares will yield the same polynomial $g(\cdot)$ that was used to generate the shares, as $D^{\prime}$ point values determine a unique polynomial of degree at most $D^{\prime}-1$. Hence, any selection of $D^{\prime}$ shares will issue the secret $s$. On the other hand, any selection of $D^{\prime}-1$ shares reveals nothing about the secret $s$ (in the sense that every value in the field remains equally probable to be the secret $s$).
2.2 Distributed tallying and secure computations over secret shares
@@
Our protocol involves a set of parties, $\{T_{1},\ldots,T_{D}\}$, that are called talliers. In the protocol, each voter creates shares of his [^2] For the sake of simplicity, we keep referring to parties by the pronoun “he”. In our context, those parties may be voters, who are humans of any gender, or talliers, that are typically (genderless) servers. private ballot and distributes them to the $D$ talliers. Since those ballots are matrices over ${\bf Z}_{p}$, as we discuss in Section 3, the secret sharing is carried out for each matrix entry independently.
@@
We will use Shamir’s $D^{\prime}$-out-of-$D$ secret sharing with $D^{\prime}:=\lfloor(D+1)/2\rfloor$ (namely, the value $(D+1)/2$ rounded down to the nearest integer). With that setting, at least half of the talliers would need to collaborate in order to recover the secret ballots. If the set of talliers is trusted to have an honest majority, then such a betrayal scenario is impossible. Namely, if more than half of the talliers are honest, in the sense that they would not attempt cheating, then even if all other dishonest talliers collude in attempt to recover the secret ballots from the shares that they hold, they will not be able to extract any information on the ballots. Higher values of $D$ (and consequently of $D^{\prime}:=\lfloor(D+1)/2\rfloor$) will imply greater security against coalitions of corrupted talliers, but they will also imply higher costs.
@@
The first task of the talliers upon receiving the shares of a voter’s ballot matrix is to verify that those shares correspond to a legal ballot (see Section 3.2 where we explain what constitutes a legal ballot). Then, at the end of the election period, the talliers need to compute the identity of the winning candidates, as dictated by the rule. These tasks would be easy if the talliers could use their shares in order to recover the ballots. However, they must not do so, in order to protect the voters’ privacy. Instead, they must perform these computations on the distributed shares, without revealing the shares to each other. As we shall see later on, these computations boil down to the four specific tasks that we proceed to describe.
@@
Let $x_{1},\ldots,x_{L}$ be $L$ secrets in the underlying field ${\bf Z}_{p}$ ($L$ is any integer) and assume that the talliers hold $D^{\prime}$-out-of-$D$ shares of each of those secrets. Then the talliers can perform the following computational tasks without recovering the secrets $x_{1},\ldots,x_{L}$ or learning anything about them:
@@
    Evaluating arithmetic functions. If $y=f(x_{1},\ldots,x_{L})$, where $f$ is some public arithmetic function, compute $D^{\prime}$-out-of-$D$ shares of $y$.
@@
    Secure comparisons. If $y=1_{x_{1}<x_{2}}$ is the bit that equals 1 if $x_{1}<x_{2}$ and 0 otherwise333 For any predicate $\Pi$, we denote by $1_{\Pi}$ the bit that equals 1 if and only if (iff hereinafter) the predicate $\Pi$ holds., compute $D^{\prime}$-out-of-$D$ shares of $y$.
@@
    Testing positivity. If $p>2N$ (where $N$ is the number of voters) and $x_{1}\in[-N,N]$, compute $D^{\prime}$-out-of-$D$ shares of the bit $1_{x_{1}>0}$.
@@
    Testing equality to zero. If $p>2N$ and $x_{1}\in[-N,N]$, compute $D^{\prime}$-out-of-$D$ shares of the bit $1_{x_{1}=0}$.
All those computational tasks can be carried out by secure multiparty computation (MPC) [51] sub-protocols. We treat those computations as black boxes. Namely, it is sufficient to understand what are their inputs and outputs, as described above, but there is no need to understand their internal operation. However, interested readers may refer to A for more information on those computations.
3 A secure order-based voting protocol